# FlexiSCD: Flexible Surface Code Deformer for Dynamic Defects

### v2

@@ VIII Related Work
(1) Hardware-level solutions. Dynamic defects can be mitigated at the hardware level to reduce their incidence rate. Physical techniques such as shielding [24, 30], phonon downconversion [57, 58], and metallic covers [59] can help mitigate the burst of quasiparticles induced by radiation materials and cosmic rays, which is a main source of qubit decoherence. Various mitigation methods on chips [21, 60] are also proposed. However, these hardware-level solutions alone are often insufficient to reduce the logical error rate to a level where meaningful quantum computations, which can take several hours [50], can be effectively executed. Although preventing these defects at the hardware level is challenging, specialized hardware detectors [31, 32] can often locate them accurately at runtime with statistical methods.
(2) Software-level solution. To address dynamic defects that involve varying error rates, methods like [2, 61, 62, 63] estimate modified error rates based on historical syndrome data and then adjust the error rates in the decoders such as MWPM  [47, 64, 65, 66, 67, 68]. Various error mitigation approaches [69, 70, 71] based on statistical methods are also proposed. However, these methods becomes ineffective when the physical error rate is considerably high or the errors can possible propagate, but these are often the case with dynamic defects, where defected qubits can have error rates near 50% and affect the integrity of adjacent qubits [22, 25].
(3) QEC-level solution. Theoretical frameworks have been established for converting one stabilizer code to another [72, 73, 74, 75, 48, 76] by modifying the code through a series of measurements. However, these methods either potentially modify the encoded information [75, 48], or ignore the hardware constraints on qubit connectivity [76], limiting their code conversion procedure for practical use. On the other hand, approaches such as [2, 77, 1] design new codes based on the structure of the underlying hardware. However, these methods are inefficient in utilizing the remaining intact qubits, do not detail the specific deformation procedure, and overlook the feasibility of communication between logical qubits.
@@ IX Conclusion
In this work, we introduce a novel deformation framework FlexiSCD, which expands the existing surface code instruction set by integrating adaptive defect mitigation functionalities. Based on four meticulously designed deformation instructions, FlexiSCD creates a larger design space than previous work, enabling optimized deformation processes tailored to specific defects. When combined with a new adaptive code layout, FlexiSCD significantly enhances qubit communication efficiency and reduces logical error rates with only mild additional qubit resource. Comprehensive evaluations show that FlexiSCD produces deformed codes with better QEC capabilities and more efficient qubit communication. FlexiSCD serves as a versatile tool for effectively addressing defects in surface codes across various quantum platforms.

### v3

@@ VIII Related Work
(1) Hardware-level solutions. Dynamic defects can be mitigated at the hardware level to reduce their incidence rate. Physical techniques such as shielding [24, 30], phonon downconversion [67, 68], and metallic covers[69] can help mitigate the burst of quasiparticles induced by radiation materials and cosmic rays, which is a main source of qubit decoherence. Various mitigation methods on chips [21, 70, 71] are also proposed. However, although these hardware-level solutions can improve the defect rate of physical qubit, it is often insufficient for meaningful quantum computations, which can take several hours [60], to be effectively executed at low retry risk. Although preventing these defects at the hardware level is challenging, specialized hardware detectors [31, 32] can often locate them accurately at runtime with statistical methods.
(2) Software-level solution. To address dynamic defects that involve varying error rates, methods like [2, 72, 73, 74] estimate modified error rates based on historical syndrome data and then adjust the error rates in the decoders such as MWPM [47, 75, 76, 48, 77, 78]. Various error mitigation approaches [79, 80, 81] based on statistical methods are also proposed. However, these methods becomes ineffective when the physical error rate is considerably high or the errors can propagate, but these are often the case with dynamic defects, where defected qubits can have error rates near 50% and affect the integrity of adjacent qubits [22, 25].
(3) QEC-level solution. Theoretical frameworks have been established for converting one stabilizer code to another [82, 83, 84, 52, 57] by modifying the code through a series of measurements. However, these methods either potentially modify the encoded information [84, 52], or ignore the hardware constraints on qubit connectivity [57], limiting their code conversion procedure for practical use. On the other hand, approaches such as [2, 50, 1] design new codes based on the structure of the underlying hardware. However, these methods are inefficient in utilizing the remaining intact qubits, do not detail the specific deformation procedure, and overlook the feasibility of communication between logical qubits.
@@ IX Conclusion
We introduce Surf-Deformer, a novel deformation framework that expands the surface code instruction set to integrate adaptive defect mitigation functionalities. With four meticulously designed deformation instructions, Surf-Deformer creates a larger design space, enabling optimized deformation processes tailored to specific defects. Combined with a new adaptive code layout, Surf-Deformer enhances qubit communication efficiency and reduces logical error rates with minimal additional qubit resources. Comprehensive evaluations show that Surf-Deformer produces deformed codes with superior QEC capabilities and more efficient qubit communication. Surf-Deformer is a versatile tool for effectively addressing defects in surface codes across various quantum platforms.
Acknowledgment
We thank the anonymous reviewers for their constructive feedback and AWS Cloud Credit for Research. This work is supported in part by Robert N.Noyce Trust, NSF 2048144, NSF 2422169, NSF 2427109. This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center. This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725. The Pacific Northwest National Laboratory is operated by Battelle for the U.S. Department of Energy under Contract DE-AC05-76RL01830.
Appendices
IX-A Proofs for the logical state preservation
In this appendix, we prove that the gauge transformation does not alter the logical state stored in the subsystem stabilizer code. Based on the stabilizer formalism [53], a $[n,k]$ subsystem stabilizer code $\mathbb{C}$, which encodes $n$ physical qubits into $k$ logical qubits, can always be represented as:
$s_{1},~{}\dots,~{}s_{n-k-l},~{}\begin{bmatrix}\bar{X}_{1}\\ \bar{Z}_{1}\end{bmatrix},~{}\dots,~{}\begin{bmatrix}\bar{X}_{k}\\ \bar{Z}_{k}\end{bmatrix},~{}\begin{bmatrix}\bar{X}_{k+1}\\ \bar{Z}_{k+1}\end{bmatrix},~{}\dots,~{}\begin{bmatrix}\bar{X}_{k+l}\\ \bar{Z}_{k+l}\end{bmatrix}$
where $s$, $\bar{X}$, and $\bar{Z}$ are length-$n$ Pauli strings representing the operators acting on the $n$ physical qubits. The operators $s_{1},\dots,s_{n-k-l}$ are the stabilizer generators of the code $\mathbb{C}$. $\bar{X}_{i}$ and $\bar{Z}_{i}$ represent the logical $X$ and $Z$ operators for the $i$-th logical qubit. Similarly, $\bar{X}_{k+i}$ and $\bar{Z}_{k+i}$ correspond to the logical $X$ and $Z$ operators for the $i$-th gauge qubit, indicating that this degree of freedom is not used to store information. We refer to this as the generator representation of code $\mathbb{C}$. For example, the original surface code has $k=1$ and $l=0$. From the theory of stabilizer formalism, we have:
Theorem 1.
A generator representation is valid for a code $\mathbb{C}$ if and only if it satisfies the following conditions:
(1) All operators are independent, meaning that the product of any subset of the operator set $\{s\}\cup\{\bar{X}\}\cup\{\bar{Z}\}$ does not equal to the identity operator $\mathbb{I}$.
(2) The two logical operators on a single logical qubit or gauge qubit must anti-commute: $\bar{X}_{i}\bar{Z}_{i}=-\bar{Z}_{i}\bar{X}_{i}$.
(3) Apart from the operator pairs in (2), any two operators in $\{s\}\cup\{\bar{X}\}\cup\{\bar{Z}\}$ must commute.
Without loss of generality, let us consider the case $k=1$ in the following part, which corresponds to the surface code setting used in our Surf-Deformer. The case for $k>1$ can be proven in a similar manner. The generator representation of a $k=1$ code $\mathbb{C}$ is:
$s_{1},~{}\dots,~{}s_{n-l-1},~{}\begin{bmatrix}\bar{X}_{L}\\ \bar{Z}_{L}\end{bmatrix},\begin{bmatrix}\bar{X}_{1}\\ \bar{Z}_{1}\end{bmatrix},~{}\dots,~{}\begin{bmatrix}\bar{X}_{l}\\ \bar{Z}_{l}\end{bmatrix}$
For any code state $|\psi\rangle$ in the code space $\mathbb{C}$, it can always be represent as:
$|\psi\rangle=\alpha|\psi_{0}\rangle+\beta|\psi_{1}\rangle$
where $|\psi_{0}\rangle$ and $|\psi_{1}\rangle$ are the logical zero and one state identified by $\bar{Z}_{L}$:
$\bar{Z}_{L}|\psi_{0}\rangle=|\psi_{0}\rangle,~{}~{}~{}\bar{Z}_{L}|\psi_{1}\rangle=-|\psi_{1}\rangle$
Then we have the definition of logical state preservation:
Definition 2.
When a code $\mathbb{C}$ is deformed to $\mathbb{C^{\prime}}$, the deformation process is said to be logical-state-preserving on $\bar{Z}_{L}$ if and only if, for any code state $|\psi\rangle$ in $\mathbb{C}$, the resulting state after deformation remains $|\psi^{\prime}\rangle=\alpha|\psi_{0}^{\prime}\rangle+\beta|\psi_{1}^{\prime}\rangle$, where $|\psi_{0}^{\prime}\rangle$ and $|\psi_{1}^{\prime}\rangle$ are the logical zero and one states identified by the new logical operator $\bar{Z^{\prime}}_{L}$ of the code $\mathbb{C^{\prime}}$.
Definition 3.
When a code $\mathbb{C}$ is deformed to $\mathbb{C^{\prime}}$, the deformation process is said to be logical-state-preserving if and only if it is logical-state-preserving on both $\bar{X}_{L}$ and $\bar{Z}_{L}$.
It is important to note that the generator representation of a code $\mathbb{C}$ is not identical to the stabilizers and gauge operators measured in practice. The set of operators actually measured by the circuit can be expressed as $Meas=Stab\cup Gauge$. The stabilizer set $Stab$ is measured in every cycle, while
To determine whether a code $\mathbb{C}$ can get all necessary error syndromes by measuring $Meas$, we define the validity of $Meas$:
Definition 4.
A operator set Meas is valid for a code $\mathbb{C}$ if and only if it satisfies the following conditions:
(1) $Stab\subseteq\langle s_{1},~{}\dots,~{}s_{n-l-1}\rangle$
(2) $Gauge\subseteq\langle s_{1},\dots,s_{n-l-1},\bar{X}_{1},\bar{Z}_{1},\dots,\bar{X}_{l},\bar{Z}_{l}\rangle\backslash\langle s_{1},\dots,s_{n-l-1}\rangle$
(3) $\{s_{1},~{}\dots,~{}s_{n-k-1}\}\subseteq\langle Meas\rangle$
For example, a trivial measurement set for the code $\mathbb{C}$ is $Stab=\{s_{1},\dots,s_{n-k-1}\}$ and $Gauge=\emptyset$. Conditions 1 and 2 ensure that the measured stabilizers and gauge operators are generated from sets that exclude the logical operators $X_{L}$ and $Z_{L}$ on the logical qubit. As a result, measuring the operators in $Stab$ and $Gauge$ will not affect the logical state within the valid code space of $\mathbb{C}$, which is a fundamental property of the subsystem stabilizer code [53].
Condition 3 ensures that the error syndrome for each stabilizer generator $s_{i}$ can be determined from the measurement results of $Stab$ and $Gauge$. For example, in fig. 6(a), the super-stabilizers $s_{1}s_{2}$ and $g_{1}g_{2}$ are two stabilizer generators in the deformed code $\mathbb{C^{\prime}}$, and their results can be inferred from $Gauge=\{s_{1}^{\prime},s_{2}^{\prime},g_{1}^{\prime},g_{2}^{\prime}\}$, as $s_{1}s_{2}=s_{1}^{\prime}s_{2}^{\prime}$ and $g_{1}g_{2}=g_{1}^{\prime}g_{2}^{\prime}$.
In the following, we discuss the four types of gauge transformations formalized in II and demonstrate how each preserves the logical state:
Stabilizer-to-Stabilizer (S2S) Transformation: The S2S transformation identifies two stabilizers, $\hat{s_{i}},\hat{s_{j}}\in Stab$, and either replaces one of them with $\hat{s_{i}}\hat{s_{j}}$ or adds $\hat{s_{i}}\hat{s_{j}}$ to the $Stab$. This transformation does not alter the generator representation of the code $\mathbb{C}$. The updated measurement set, $Meas^{\prime}$, remains valid for the code $\mathbb{C}$, since the transformation does not change the generated group, which means $\langle Meas\rangle=\langle Meas^{\prime}\rangle$. Therefore, measuring the new $Meas^{\prime}$ will not disturb the logical state.
Gauge-to-Gauge (G2G) Transformation: The G2G transformation identifies a gauge operator, $\hat{g}\in Gauge$, and a measurement operator, $\hat{m}\in Meas$. If $\hat{g}\hat{m}\notin\langle s_{1},\dots,s_{n-l-1}\rangle$, the transformation either replaces $\hat{g}$ with $\hat{g}\hat{m}$ or adds $\hat{g}\hat{m}$ to $Gauge$. Like S2S transformation, G2G transformation will not alter the generator representation of the code $\mathbb{C}$ and disturb the logical state.
Stabilizer-to-Gauge (S2G) Transformation: The S2G transformation identifies a new operator $\hat{g}$ and all the stabilizers that anti-commute with it:
$Anti=\{\hat{s_{i}}\in Stab~{}|~{}\hat{g}\hat{s_{i}}=-\hat{s_{i}}\hat{g}\}$
If $Anti\neq\emptyset$, this transformation adds $\hat{g}$ to $Gauge$ and moves elements of $Anti$ from $Stab$ to $Gauge$. It follows that $Meas^{\prime}=Meas\cup\{g\}$, meaning the only practical change is the measurement of the new gauge operator $\hat{g}$.
To illustrate how the code $\mathbb{C}$ is deformed, let us first provide an equivalent generator representation for the code $\mathbb{C}$.
Theorem 5.
For any code $\mathbb{C}$ and any new operator $g$, if there exists a stabilizer generator $s_{k}$ that anti-commutes with $g$, there is an alternative generator representation of the code $\mathbb{C}$:
$s_{1}^{\prime},~{}\dots,~{}s_{k},~{}\dots,~{}s_{n-l-1}^{\prime},~{}\begin{bmatrix}\bar{X}_{L}^{\prime}\\ \bar{Z}_{L}^{\prime}\end{bmatrix},\begin{bmatrix}\bar{X}_{1}^{\prime}\\ \bar{Z}_{1}^{\prime}\end{bmatrix},~{}\dots,~{}\begin{bmatrix}\bar{X}_{l}^{\prime}\\ \bar{Z}_{l}^{\prime}\end{bmatrix}$
where $g$ commutes with all operators in the set $\{s^{\prime}\}\cup\{\bar{X^{\prime}}\}\cup\{\bar{Z^{\prime}}\}$.
Proof: If $s_{i}\neq s_{k}$ anti-commutes with $g$, define $s_{i}^{\prime}=s_{k}s_{i}$. Similarly, if $\bar{X_{i}}$ (or $\bar{Z_{i}}$) anti-commutes with $g$, define $\bar{X_{i}}^{\prime}=s_{k}\bar{X_{i}}$ (or $Z_{i}^{\prime}=s_{k}\bar{Z_{i}}$). Since this replacement does not alter the generated stabilizer group, logical operator group, or gauge operator group, it provides an alternative generator representation for the code $\mathbb{C}$, where only $s_{k}$ anti-commutes with $g$.
For $Anti\neq\emptyset$, we can always find a stabilizer generator $s_{i}$ that anti-commutes with the new gauge operator $\hat{g}$. This is because any stabilizer $\hat{s}$ in $Anti$ can be expressed as a product of stabilizer generators, and at least one of these generators must anti-commute with $\hat{g}$.
Theorem 6.
While preserving the logical state of code $\mathbb{C}$, the S2G transformation with new gauge operator $g$ deforms it into a new code $\mathbb{C^{\prime}}$:
$s_{1}^{\prime},~{}\dots,~{}\begin{bmatrix}s_{k}\\ g\end{bmatrix},\dots,~{}s_{n-l-1}^{\prime},~{}\begin{bmatrix}\bar{X}_{L}^{\prime}\\ \bar{Z}_{L}^{\prime}\end{bmatrix},\begin{bmatrix}\bar{X}_{1}^{\prime}\\ \bar{Z}_{1}^{\prime}\end{bmatrix},~{}\dots,~{}\begin{bmatrix}\bar{X}_{l}^{\prime}\\ \bar{Z}_{l}^{\prime}\end{bmatrix}$
which only replace stabilizer generator $s_{i}$ with a pair of gauge operators $(s_{k},~{}g)$.
Proof: For any code state $|\psi\rangle=\alpha|\psi_{0}\rangle+\beta|\psi_{1}\rangle$, we can denote:
$|\psi_{0}\rangle$ $=\frac{|\psi_{0}\rangle+g|\psi_{0}\rangle}{2}+\frac{|\psi_{0}\rangle-g|\psi_{0}\rangle}{2}:=\frac{|P_{0}\rangle}{\sqrt{2}}+\frac{|N_{0}\rangle}{\sqrt{2}}$
$|\psi_{1}\rangle$ $=\frac{|\psi_{1}\rangle+g|\psi_{1}\rangle}{2}+\frac{|\psi_{1}\rangle-g|\psi_{1}\rangle}{2}:=\frac{|P_{1}\rangle}{\sqrt{2}}+\frac{|N_{1}\rangle}{\sqrt{2}}$
since $s_{k}|P_{0/1}\rangle=|N_{0/1}\rangle$. It is easy to see that
$g|P_{0/1}\rangle$ $=|P_{0/1}\rangle,$
$g|N_{0/1}\rangle$ $=-|N_{0/1}\rangle.$
Thus, after measurement of $g$, the resulting state will be $|\psi^{\prime}\rangle=\alpha|P_{0}\rangle+\beta|P_{1}\rangle$ if the measurement result is 0, or $|\psi^{\prime}\rangle=\alpha|N_{0}\rangle+\beta|N_{1}\rangle$ if the measurement result is 1.
Since $g$ commutes with the logical operator $\bar{Z_{L}}^{\prime}$, the states $|P_{0}\rangle$ and $|N_{0}\rangle$ correspond to the logical zero state identified by $\bar{Z_{L}}^{\prime}$, while $|P_{1}\rangle$ and $|N_{1}\rangle$ correspond to the logical one state.
The proof for $X_{L}$ can be derived in a similar manner. Therefore, the process of S2G transformation is logical-state-preserving.
Gauge-to-Stabilizer (G2S) Transformation: The G2G transformation identifies a gauge operator $\hat{g}\in Gauge$ and all gauge operators that anti-commute with it:
$Anti=\{\hat{g_{i}}\in Gauge~{}|~{}\hat{g}\hat{g_{i}}=-\hat{g_{i}}\hat{g}\}$
If $|Anti|>1$, perform additional G2G transformations until $|Anti|=1$. This ensures that the measurement set $Meas^{\prime}$ remains valid for the deformed code $\mathbb{C^{\prime}}$. Finally, this transformation adds $\hat{g}$ to the stabilizer set $Stab$ and removes the elements of $Anti$ from the gauge set. The changes in the generator representation can be considered the reverse process of the S2G transformation. In practice, we only measure $\hat{g}$ and apply the $s_{k}$ operation if the result is 1. This does not affect the logical state, as $s_{k}|P_{0/1}\rangle=|N_{0/1}\rangle$.